OpenStudy (anonymous):

The width of a rectangle is 1ft less than the length. The area is 20ft^2. Find the length and the width.

OpenStudy (p0sitr0n):

length = x width = x-1 A = L * W A = x * (x-1) A = 20 20 = x^2 -x x^2-x-20 =0 x = 5 width =4